### v4

# Residual-based Chebyshev filtered subspace iteration for sparse Hermitian eigenvalue problems tolerant to inexact matrix-vector products with applications to Kohn-Sham density functional theory

Abstract
Chebyshev Filtered Subspace Iteration (ChFSI) has emerged as a robust alternative to Krylov eigensolvers for extracting a small subset of extremal eigenpairs from large sparse matrices, particularly in situations where these eigenpairs must be computed repeatedly as the system matrix evolves within an outer iteration. Such scenarios arise prominently in computational physics, including ab initio materials modelling using density functional theory (DFT), as well as in certain stability and coupled multiphysics analyses where eigenvalue information augments nonlinear solution procedures. In this work, we propose R-ChFSI, a residual-based reformulation of ChFSI designed to exhibit strong convergence properties even when the matrix–vector products are computed inexactly. By recasting the Chebyshev polynomial recurrence in terms of residuals rather than direct eigenvector updates, R-ChFSI effectively suppresses errors introduced by approximate operator applications and preserves the convergence behaviour of the ChFSI scheme without approximations. We further derive convergence guarantees under matrix–vector product approximations, providing a rigorous foundation for the method in large-scale eigenvalue computations. The tolerance of R-ChFSI to inexact matrix-vector products enables an efficient treatment of generalized Hermitian definite eigenproblems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, where exact factorizations or high-accuracy iterative solves for evaluating $\textbf{B}^{-1}$ are often prohibitively expensive. By allowing the use of inexpensive approximate inverses of B, the method substantially reduces computational costs while maintaining robust convergence. Moreover, R-ChFSI naturally accommodates low-precision arithmetic for both standard and generalized eigenproblems, making it well-suited for modern hardware accelerators optimised for mixed-precision computation. To demonstrate the effectiveness of the approach, extensive numerical experiments are conducted on finite-element discretized eigenproblems with millions of degrees of freedom, solving for thousands of eigenpairs arising in ab initio material modelling using Kohn-Sham density functional theory. For generalized eigenproblems employing approximate $\textbf{B}^{-1}$, R-ChFSI achieves desired residual norms orders of magnitude smaller than those obtained with standard ChFSI. In addition, R-ChFSI reliably reaches target residual tolerances (e.g., 10-8) even with FP32 and TF32 arithmetic, significantly outperforming standard ChFSI in similar settings. By reducing dependence on high-precision arithmetic and enabling efficient use of modern accelerator hardware, R-ChFSI provides a scalable and computationally efficient eigensolver for large sparse eigenproblems central to ab initio electronic structure calculations, multiphysics simulations, and large-scale engineering analysis.
keywords:
Sparse Hermitian generalized eigenproblems, Chebyshev filtered subspace iteration, finite-element basis, Kohn-Sham density functional theory
journal: Nuclear Physics B
[…]
[inst1]organization=Department of Computational and Data Sciences, Indian Institute of Science,addressline=CV Raman Road, city=Bengaluru, postcode=560012, state=Karnataka, country=India
1 Introduction
Large sparse Hermitian eigenvalue problems play a central role in many areas of computational physics. They arise from the discretization of fundamental governing equations in quantum mechanics, electronic structure theory, fluid dynamics, plasma physics, wave propagation, electromagnetics, and elastodynamics. In such applications, the extremal eigenmodes often correspond to physically meaningful states—ground and low-lying excited states in quantum systems, dominant instability modes in fluid or plasma flows, or fundamental vibration frequencies in elastic media—which makes the efficient computation of only a small subset of the spectrum particularly important. The solution strategies for these eigenproblems often rely on fully iterative methods based on iterative orthogonal projection approaches. In these approaches, the large sparse matrix is orthogonally projected onto a carefully constructed smaller subspace rich in the wanted eigenvectors (Rayleigh-Ritz step), followed by subspace diagonalisation of the projected matrix and a subspace rotation step to recover the desired orthogonal eigenvector estimates of the original sparse Hermitian matrix. Popular iterative approaches include Davidson [crouzeixDavidsonMethod1994, davidsonIterativeCalculationFew1975], Generalized-Davidson [morganGeneralizationsDavidsonsMethod1986], Jacobi-Davidson [hochstenbachJacobiDavidsonMethod2006], Chebyshev-filtered subspace iteration (ChFSI) approach [zhouSelfconsistentfieldCalculationsUsing2006], LOBPCG [knyazevOptimalPreconditionedEigensolver2001], and PPCG [vecharynskiProjectedPreconditionedConjugate2015]. Another key class of iterative techniques is based on Krylov subspace methods, namely the Arnoldi method [arnoldiPrincipleMinimizedIterations1951], Lanczos methods [lanczosIterationMethodSolution1950] and their important variants, including implicit restart Arnoldi methods [sorensenImplicitApplicationPolynomial1992], Krylov-Schur method [stewartKrylovSchurAlgorithmLarge2002] and block-Krylov methods [saadNumericalMethodsLarge2011]. Our focus in this work involves the solution of large, sparse Hermitian eigenproblems using the ChFSI approach.
Chebyshev filtered subspace iteration (ChFSI) has emerged as a robust alternative to Krylov subspace eigensolvers for extracting a small subset of extremal eigenpairs from large, sparse Hermitian matrices—particularly in scenarios where these eigenpairs must be computed repeatedly as the system matrix evolves within an outer iteration. A prominent example is the nonlinear eigenvalue problem that arises in Kohn-Sham density functional theory (DFT) [kohnSelfconsistentEquationsIncluding1965], where one is interested in the lowest eigenpairs of a Hamiltonian that depends nonlinearly on the eigenvectors. This problem is typically solved through a Self-Consistent Field (SCF) procedure, in which the Hamiltonian is updated at each iteration and a small set of extremal eigenpairs must be recomputed repeatedly as the system approaches self-consistency. ChFSI relies on a Chebyshev polynomial filtering procedure that constructs a subspace rich in the desired eigenvectors, exploiting the fast growth property of Chebyshev polynomials outside the interval [-1,1], followed by a Rayleigh-Ritz step of projection and subspace diagonalization. Approaches based on ChFSI have become quite popular recently to solve both standard and generalized eigenproblems arising in DFT [lohChebyshevFilteringLanczos1984, zhouParallelSelfconsistentfieldCalculations2006, motamarriHigherorderAdaptiveFiniteelement2013, levittParallelEigensolversPlanewave2015, wuAdvancingDistributedMultiGPU2023]. Although variants of ChFSI approaches combined with the Davidson method [zhouChebyshevDavidsonAlgorithm2007, miaoChebyshevDavidsonMethod2020, miaoFlexibleBlockChebyshevDavidson2023, wangNewSubspaceIteration2025] have been proposed, the ChFSI approach remains the preferred choice in electronic structure codes based on DFT for solving the underlying nonlinear eigenvalue problem [dasDFTFE10Massively2022, ghoshSPARCAccurateEfficient2017a, zhouChebyshevfilteredSubspaceIteration2014, fengMassivelyParallelImplementation2024, banerjeeChebyshevPolynomialFiltered2016] due to its scalability, ability to accommodate evolving subspaces and computational efficiency [dasFastScalableAccurate2019, dasLargeScaleMaterialsModeling2023]. Additionally, ChFSI is memory efficient compared to Davidson-type approaches and further offers an attractive alternative to preconditioned-type conjugate gradient approaches for problems, especially for problems where good preconditioners are unavailable. The current work fills an important gap in the ChFSI approach by developing a residual-based approach to ChFSI that is tolerant to inexact matrix-vector products during the subspace construction step. Consequently, we argue that the method is well-suited for solving generalized eigenproblems efficiently and additionally low-precision arithmetic can be leveraged in the light of recent changes in modern heterogeneous computing architectures, significantly improving computational efficiency.
Modern hardware architectures have undergone substantial modifications in recent years due to the computational requirements of machine learning (ML) and artificial intelligence (AI) training. Owing to their high computational demands, these domains have gravitated towards the use of low-precision arithmetic for training and inferencing. In response to this demand, hardware manufacturers have been enhancing support for low-precision floating-point formats, such as tensorfloat32 and bfloat16, enabling significantly faster throughput and substantial performance improvements [^1] NVIDIA’s Blackwell GPUs, designed for AI/ML applications, demonstrate a notable decrease in peak double-precision (FP64) floating-point performance [HGXAISupercomputing] compared to the previous Hopper architecture.. These architectural changes underscore the necessity to modify scientific computing algorithms to efficiently utilize low-precision processes without compromising accuracy [dongarraHardwareTrendsImpacting2024, kashiMixedprecisionNumericsScientific2025]. For sparse eigensolvers, the ideas of mixed precision preconditioning for the LOBPCG eigensolver [kressnerMixedPrecisionLOBPCG2023] and mixed precision orthogonalization and Rayleigh-Ritz have been explored for the LOBPCG and ChFSI eigensolvers [kressnerMixedPrecisionLOBPCG2023, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020]. We note that in these works the evaluation of matrix-vector products involved in the subspace construction still need to be performed in the higher precision arithmetic. The development of robust iterative eigensolver algorithms to accommodate low precision arithmetic in matrix-vector multiplications becomes particularly important in light of the evolving hardware landscape.
In this work, we introduce a residual-based reformulation of the Chebyshev filtered subspace iteration (ChFSI) method [saadChebyshevAccelerationTechniques1984, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, zhouChebyshevfilteredSubspaceIteration2014], referred to as the R-ChFSI method, for solving large-scale sparse Hermitian eigenvalue problems. The key novelty of R-ChFSI lies in its ability to accommodate inexact matrix-vector products in the subspace construction step while preserving convergence properties. Most importantly, for generalized eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$, we demonstrate that the method naturally admits low-cost approximations of $\textbf{B}^{-1}$, allowing ChFSI to act on $\textbf{B}^{-1}\textbf{A}$ and solve large sparse generalized eigenproblems without degrading convergence, thereby avoiding expensive matrix factorizations and iterative solvers that are otherwise typically required for such problems [levittParallelEigensolversPlanewave2015]. The ability of R-ChFSI to be robust to approximations in matrix-vector multiplications makes the proposed method particularly relevant for modern hardware architectures that favor low-precision arithmetic, significantly improving computational efficiency. The core of our approach is a reformulated recurrence relation that modifies the standard ChFSI update step to operate on residuals rather than the guess of the eigenvectors. We then provide a mathematical justification demonstrating that this reformulation reduces the numerical error in the Chebyshev filtered subspace construction while employing inexact matrix-vector products compared to the traditional ChFSI recurrence relation. Our analysis establishes that the proposed R-ChFSI method effectively controls error propagation in the Chebyshev filtering process, resulting in more reliable convergence.
Through extensive benchmarking on large-scale sparse generalized eigenvalue problems arising in ab initio calculations using finite-element (FE) discretization of DFT, where we employ a diagonal approximation to the FE basis overlap matrix inverse, R-ChFSI attains a residual tolerance that is orders of magnitude lower than standard ChFSI, demonstrating superior robustness to approximation errors. The improved convergence for generalized eigenvalue problems is particularly noteworthy, as it highlights the ability of R-ChFSI to accommodate approximate inverses while maintaining accuracy. In contrast, standard ChFSI is more sensitive to such approximations, often leading to a loss in accuracy or requiring significantly higher computational costs to achieve similar residual tolerances. This advantage of R-ChFSI is especially relevant in DFT calculations, where generalized eigenvalue problems frequently arise [levittParallelEigensolversPlanewave2015, ramakrishnanFastScalableFiniteelement2025, dasDFTFE10Massively2022], and high computational costs in the subspace construction constitute a significant bottleneck. Furthermore, R-ChFSI enables the use of lower-precision arithmetic in the filtering step, further enhancing the performance gains. In our benchmark studies, using TF32 arithmetic in Intel Data Center Max GPU accelerators deployed on the supercomuting system Aurora, we have obtained speedups of up to 2.3x for the filtering step and 1.9x for the complete eigensolve to reach the desired tolerance. Further, we have also employed reduced precision, BF16, for the nearest neighbor MPI communication arising in sparse matrix-vector multiplications to enhance the performance and we obtain speedups of up to 2.7x for the filtering step and 2.1x for the complete eigensolve to reach the desired tolerance. By reducing dependence on exact matrix factorizations and high-precision arithmetic, R-ChFSI enables efficient exploitation of modern hardware accelerators, making large-scale eigenvalue computations more feasible in high-performance computing environments.
The remainder of this article is organized as follows: Section 2 outlines the key steps of the Chebyshev filtered subspace iteration procedure and subsequently analyzes the convergence properties of ChFSI in terms of how the maximum principal angle between the current subspace and the target eigenspace evolves during the iterations. Section 3 begins by analyzing the convergence of ChFSI when the subspace is constructed approximately due to inexact matrix-vector products. Subsequently, this section introduces the proposed residual-based Chebsyshev filtered subspace iteration method (R-ChFSI) for generalized eigenvalue problems. This section discusses the convergence of R-ChFSI with approximate matrix products and demonstrates mathematically that the R-ChFSI method can converge even when the traditional ChFSI method fails. Section 4 presents a comprehensive evaluation of the proposed R-ChFSI method in terms of accuracy and computational efficiency on GPU architectures that admit low precision arithmetic. It compares R-ChFSI to the traditional ChFSI approach for solving real symmetric and complex Hermitian eigenproblems in the case of generalized eigenproblems while using inexact matrix-vector products.
2 Mathematical Background
The ChFSI [saadChebyshevAccelerationTechniques1984, saadNumericalMethodsLarge2011, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, motamarriHigherorderAdaptiveFiniteelement2013, zhouChebyshevfilteredSubspaceIteration2014] approach for solving the desired eigenpairs belongs to the category of iterative orthogonal projection methods and is one of the widely used strategies to compute the smallest $n$ eigenvalues and their corresponding eigenvectors [dektorInexactSubspaceProjection2025, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020, zhangSPARCV200Spinorbit2024, ghoshSPARCAccurateEfficient2017a, xuSPARCSimulationPackage2021, kronikPARSECPseudopotentialAlgorithm2006, liouScalableImplementationPolynomial2020, winkelmannChASE2019, michaud-riouxRESCURealSpace2016, doganSolvingElectronicStructure2023, levittParallelEigensolversPlanewave2015]. To describe the proposed eigensolver strategy based on the ChFSI approach, in the current work, we consider the Hermitian generalized eigenvalue problem of the form
$\textbf{A}\textbf{u}_{i}=\lambda_{i}\textbf{B}\textbf{u}_{i}\;\;\text{where}\;\;\textbf{A}\in\mathbb{C}^{m\times m},\;\textbf{B}\in\mathbb{C}^{m\times m}.$ (1)
where A and B are Hermitian matrices with B being a positive-definite matrix. In addition, $\lambda_{i}\in\mathbb{R}$ and $\textbf{u}_{i}\in\mathbb{C}^{m}:\forall i=1,\dots,n$ denote the eigenvalue-eigenvector pairs corresponding to the smallest $n$ eigenvalues. Equation 1 reduces to a standard eigenvalue problem when $\textbf{B}=\textbf{I}$ with I denoting the $m\times m$ identity matrix. Without loss of generality, we assume that the eigenvalues are ordered as $\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{n}\leq\lambda_{n+1}\leq\dots\leq\lambda_{m}$. For completeness and to introduce notations, we now provide a brief overview of the ChFSI algorithm, traditionally used to solve the eigenvalue problem corresponding to eq. 1.
2.1 Chebyshev filtered subspace iteration and convergence properties
ChFSI leverages the properties of Chebyshev polynomials to efficiently filter out the components of the unwanted eigenvectors (corresponding to the remaining mnm-n largest eigenvalues), thus enriching the trial subspace with the desired eigenvectors. To this end, we define the Chebyshev polynomial of degree $k$ as $T_{k}(x)$ and note that these polynomials exhibit the fastest growth in their magnitude when $\lvert x\rvert\,>1$ while remaining bounded between [1,1][-1,1] when $\lvert x\rvert\,\leq 1$. To exploit this fast growth property of $T_{k}(x)$, an affine transformation that maps the largest mnm-n eigenvalues to [1,1][-1,1] is defined, and consequently, the desired smallest $n$ eigenvalues get mapped to values lying in $(-\infty,-1]$. To this end, we define the center of the unwanted spectrum as $c=(\lambda_{n+1}+\lambda_{m})/2$ and the half-width of the unwanted spectrum as $e=(\lambda_{m}-\lambda_{n+1})/2$ and hence the required affine transformation can now be represented as $\mathcal{L}(x)=(x-c)/e$. Standard implementations of ChFSI also scale the Chebyshev polynomials to prevent overflow [saadChebyshevAccelerationTechniques1984, saadNumericalMethodsLarge2011, zhouParallelSelfconsistentfieldCalculations2006, zhouSelfconsistentfieldCalculationsUsing2006, zhouChebyshevfilteredSubspaceIteration2014, winkelmannChASE2019] and consequently the scaled and shifted Chebyshev polynomials are defined as $C_{k}(x)=T_{k}(\mathcal{L}(x))/T_{k}(\mathcal{L}(a_{low}))$ where $a_{low}\leq\lambda_{1}$. We note that the ChFSI procedure for solving the eigenproblem in eq. 1 for the case of sparse matrices [^2] For dense matrices the common practice is to employ Cholesky factorization of B to convert the generalized eigenvalue problem to the standard eigenvalue problem. is usually devised with the matrix $\textbf{H}=\textbf{B}^{-1}\textbf{A}$ that has the same eigenpairs as $\textbf{A}\textbf{u}_{j}=\lambda_{j}\textbf{B}\textbf{u}_{j}$. This procedure for computing the smallest eigenpairs of eq. 1 up to a specified tolerance $\tau$ on the eigenproblem residual norm, is summarized in algorithm 1. We note that in algorithm 1 we do not explicitly perform an orthogonalization step for the filtered subspace and instead opt to perform the Rayleigh-Ritz step for a non-orthogonal basis, which ensures that the resulting Ritz vectors are B-orthonormal. This choice was made for both ease of analysis and computational efficiency.
Algorithm 1 Subspace Iteration accelerated using Chebyshev polynomial of degree $p$ (ChFSI)
Initial Guess: Let $\textbf{X}^{(0)}=\begin{bmatrix}\textbf{x}_{1}^{(0)}&\textbf{x}_{2}^{(0)}&\dots&\textbf{x}_{n}^{(0)}\end{bmatrix}$ be the initial guess of the eigenvectors ($\{\textbf{u}_{j}\}$).
while $r_{j}^{(i+1)}=\lVert\textbf{A}\textbf{x}^{(i+1)}_{j}-\epsilon_{j}^{(i+1)}\textbf{B}\textbf{x}^{(i+1)}_{j}\rVert\,\geq\tau$ do
  Chebyshev Filtered Subspace Construction: Construct $\textbf{Y}_{p}^{(i)}=C_{p}(\textbf{H})\textbf{X}^{(i)}$ using the recurrence:
$\textbf{Y}_{k+1}^{(i)}=\frac{2\sigma_{k+1}}{e}\textbf{H}\textbf{Y}_{k}^{(i)}-\frac{2\sigma_{k+1}c}{e}\textbf{Y}_{k}^{(i)}-\sigma_{k}\sigma_{k+1}\textbf{Y}_{k-1}^{(i)}$ (2)
where $\textbf{Y}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\textbf{Y}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$ and $\sigma_{k+1}=1/\left(\sigma_{k}-\frac{2}{\sigma_{1}}\right)$ with $\sigma_{1}=e/(a_{low}-c)$. Here $\textbf{Y}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}$ for $k=0,1,\dots,p$ as described in algorithm 2
  Rayleigh-Ritz step: Solve the smaller $n\times n$ dense generalized eigenvalue problem. Denoting the transpose conjugate of a matrix by we have ${\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{A}\textbf{Y}_{p}^{(i)}\textbf{E}={\textbf{Y}_{p}^{(i)}}^{\dagger}\textbf{B}\textbf{Y}_{p}^{(i)}\textbf{E}\boldsymbol{\Lambda}$, where E is the eigenvector matrix and $\boldsymbol{\Lambda}$ is the diagonal matrix with the eigenvalues $\left\{\epsilon_{j}^{(i+1)}\right\}_{j=1}^{n}$ as its entries. The Ritz vectors given by $\textbf{X}^{(i+1)}=\textbf{Y}_{p}^{(i)}\textbf{E}$ are B-orthonormal while the Ritz-Values are given by $\boldsymbol{\Lambda}^{(i+1)}=\boldsymbol{\Lambda}$.
end while
Algorithm 2 Chebyshev filtering procedure for generalized Hermitian eigenvalue problems
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}$
TEMPORARY VARIABLES:X, Y
$e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$; $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$; $\sigma_{1}\leftarrow\sigma$; $\gamma\leftarrow\frac{2}{\sigma_{1}}$
$\textbf{X}\leftarrow\textbf{X}^{(i)}$; $\textbf{Y}\leftarrow\frac{\sigma_{1}}{e}(\textbf{H}-c\textbf{I})\textbf{X}^{(i)}$
for $k\leftarrow 2$ to $p$ do
  $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$
  $\textbf{X}\leftarrow\frac{2\sigma_{2}}{e}\textbf{H}\textbf{Y}-\frac{2\sigma_{2}}{e}c\textbf{Y}-\sigma\sigma_{2}\textbf{X}$
[…]
end for
return X
2.1.1 Convergence analysis of ChFSI
To ensure completeness, we now provide a mathematical justification for the convergence of ChFSI. Note that $\lVert\cdot\rVert\,$ implies the 2-norm for vectors and vector-induced matrix 2-norm for matrices (spectral norm) throughout this work. In order to analyze convergence, we first define the maximum principal angle between two subspaces [chenSpectralMethodsData2021, zhuAnglesSubspacesTheir2013, watkinsMatrixEigenvalueProblem2007].
Definition 2.1.
Let $\mathcal{X}\subset\mathbb{C}^{m\times m}$ and $\mathcal{Y}\subset\mathbb{C}^{m\times m}$ be subspaces of dimension $n$. The maximum principal angle between the two subspaces denoted by $\angle(\mathcal{X},\mathcal{Y})$ is defined as
$\sin\angle(\mathcal{X},\mathcal{Y})=\lVert\textbf{X}_{\perp}^{\dagger}\textbf{Y}\rVert\,=\lVert\textbf{Y}_{\perp}^{\dagger}\textbf{X}\rVert\,$
where $\textbf{X}\in\mathbb{C}^{m\times n}$ and $\textbf{Y}\in\mathbb{C}^{m\times n}$ are matrices whose columns form an orthonormal basis for $\mathcal{X}$ and $\mathcal{Y}$ respectively. $\textbf{X}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ and $\textbf{Y}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ are matrices whose columns form an orthonormal basis spanning the orthogonal complements of $\mathcal{X}$ and $\mathcal{Y}$ respectively.
[…]
Theorem 2.2.
For an n-dimensional space $\mathcal{S}^{(i)}$ satisfying $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ where $\mathcal{S}_{\perp}$ is the orthogonal complement of $\mathcal{S}$ and $\mathcal{S}^{(i+1)}=C_{p}(\hat{\textbf{H}})\mathcal{S}^{(i)}$,we have the following inequality
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
Proof.
Since the maximum principal angle given by definition 2.1 is invariant under unitary transformations, we have ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$ and also ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\hat{\textbf{U}}^{\dagger}\mathcal{S})}$. Since the space defined by $\hat{\textbf{U}}^{\dagger}\mathcal{S}$ is the same as $\mathcal{E}=\mathcal{R}(\textbf{E})$, we have, ${\angle(\mathcal{S}^{(i+1)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i+1)},\mathcal{E})}$ and ${\angle(\mathcal{S}^{(i)},\mathcal{S})}={\angle(\hat{\textbf{U}}^{\dagger}\mathcal{S}^{(i)},\mathcal{E})}$. Consider a partitioning of the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$ as $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible, and consequently, we can write
[…]
Subsequently from lemma A.1 we have $\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}=\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,$. Further, partitioning the diagonal matrix of eigenvalues $\boldsymbol{\Lambda}$, we have $\boldsymbol{\Lambda}=\Bigl[\begin{smallmatrix}\boldsymbol{\Lambda}_{1}&0\\ 0&\boldsymbol{\Lambda}_{2}\end{smallmatrix}\Bigr]$ where $\boldsymbol{\Lambda}_{1}$ is the $n\times n$ diagonal matrix comprising of the wanted eigenvalues and $\boldsymbol{\Lambda}_{2}$ is the $(m-n)\times(m-n)$ diagonal matrix comprising of the unwanted eigenvalues. We can now write
[…]
From lemma A.1 we now have $\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}=\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,$, allowing us to write
$\tan{\angle(\mathcal{S}^{(i+1)},\mathcal{S})}\leq\lVert C_{p}(\boldsymbol{\Lambda}_{2})\rVert\,\lVert\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,\lVert C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,=\left\lvert\frac{C_{p}(\lambda_{n+1})}{C_{p}(\lambda_{n})}\right\rvert\,\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
Theorem 2.2 demonstrates as the iterations progress, the Chebyshev filtered subspace approaches the wanted eigenspace.
3 Approximate Chebyshev Filtered Subspace Construction
We note that the computationally dominant step in Chebyshev filtered subspace construction is the evaluation of the sparse-matrix multi-vector product $\textbf{H}\textbf{Y}_{k}^{(i)}$ in eq. 2. A straightforward way to accelerate this step is to use approximations in the computation of $\textbf{H}\textbf{Y}_{k}^{(i)}$, allowing for improved efficiency. This should, in principle, allow for the use of various efficient approximate matrix multiplication techniques, including but not limited to mixed-precision arithmetic. However, we note that this requires an understanding of the convergence properties of ChFSI when such approximations are employed. We now adapt theorem 2.2 for the case where approximations are used in computing $\textbf{H}\textbf{Y}_{k}^{(i)}$.
3.1 Convergence of ChFSI with inexact subspace construction
We now define $\underline{\mathcal{S}}^{(i+1)}$ as the space spanned by the columns of $\hat{\underline{\textbf{Y}}}_{p}^{(i)}=\textbf{B}^{\frac{1}{2}}{\underline{\textbf{Y}}}_{p}^{(i)}$ with ${\underline{\textbf{Y}}}_{p}^{(i)}$ defined as $\underline{\textbf{Y}}_{p}^{(i)}=\underline{C_{p}(\textbf{H})\textbf{X}^{(i)}}$ where underline here denotes that approximations are introduced during matrix multiplications in Step 2 of algorithm 1. We note that the columns of $\hat{\textbf{X}}^{(i)}=\textbf{B}^{\frac{1}{2}}\textbf{X}^{(i)}$ form an orthonormal basis for $\mathcal{S}^{(i)}$.
Theorem 3.1.
For an n-dimensional space $\mathcal{S}^{(i)}$ satisfying $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ and $\underline{\mathcal{S}}^{(i+1)}=\mathcal{R}\left(\underline{C_{p}(\textbf{H})\textbf{X}^{(i)}}\right)$ we can write
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (3)
where $\hat{\boldsymbol{\Delta}}_{p}^{(i)}=\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}-C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}$
Proof.
Partitioning the matrix $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}$, we have $\hat{\textbf{U}}^{\dagger}\hat{\textbf{X}}^{(i)}=\Bigl[\begin{smallmatrix}\hat{\textbf{Z}}_{1}^{(i)}\\ \hat{\textbf{Z}}_{2}^{(i)}\end{smallmatrix}\Bigr]$ where $\hat{\textbf{Z}}_{1}^{(i)}=\hat{\textbf{U}}_{1}^{\dagger}\hat{\textbf{X}}^{(i)}$ and $\hat{\textbf{Z}}_{2}^{(i)}=\hat{\textbf{U}}_{2}^{\dagger}\hat{\textbf{X}}^{(i)}$. We note that the assumption $\mathcal{S}^{(i)}\cap\mathcal{S}_{\perp}=\left\{0\right\}$ ensures that $\hat{\textbf{Z}}_{1}^{(i)}$ is invertible. Consequently, we can write
$\hat{\textbf{U}}^{\dagger}\underline{\mathcal{S}}^{(i+1)}$ $=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}\underline{C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}}\right)=\mathcal{R}\left(\hat{\textbf{U}}^{\dagger}C_{p}(\hat{\textbf{H}})\hat{\textbf{X}}^{(i)}+\hat{\textbf{U}}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)$
[…]
Consequently, from lemma A.1 we can write
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $=\left\lVert\left(C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}+\hat{\textbf{U}}_{2}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)\left(C_{p}(\boldsymbol{\Lambda}_{1})\hat{\textbf{Z}}_{1}^{(i)}+\hat{\textbf{U}}_{1}^{\dagger}\hat{\boldsymbol{\Delta}}_{p}^{(i)}\right)^{-1}\right\rVert\,$
$\leq\frac{\lVert C_{p}(\boldsymbol{\Lambda}_{2})\hat{\textbf{Z}}_{2}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}{1-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,}$ (4)
Note that in the last step, we have assumed that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}C_{p}(\boldsymbol{\Lambda}_{1})^{-1}\rVert\,<1$, and using the fact that $\lVert{{}\hat{\textbf{Z}}_{1}^{(i)}}^{-1}\rVert\,=\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}$, a sufficient condition for this to be true is $\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$. Upon further simplification using sub-multiplicative property of matrix spectral norms and triangle inequality, the inequality in eq. 4 can be written as
$\tan{\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})}$ $\leq\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,}{\lvert C_{p}(\lambda_{n})\rvert\,\cos{\angle(\mathcal{S}^{(i)},\mathcal{S})}-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,}$
$=\left(\frac{\lvert C_{p}(\lambda_{n+1})\rvert\,+\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}}\right)\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
which proves the desired inequality in the theorem. ∎
Now for convergence we demand that $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ and consequently we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,\!>\!\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\;\forall i=0,\dots,\infty$ (5)
We note that if $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ remains nearly constant with iteration $i$, the right-hand side of the inequality in eq. 5 keeps increasing as we approach the exact eigenspace and beyond a certain angle $\angle(\mathcal{S}^{(i)},\mathcal{S})$ this inequality gets violated. Consequently, the angle stops decreasing and we cannot approach the exact eigenspace beyond that point. To this end, for robust convergence, we require that $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ also decreases as we approach the exact eigenspace, and we demonstrate that our proposed residual-based reformulation of Chebyshev filtered subspace iteration procedure (R-ChFSI) algorithm described subsequently accomplishes this. We further note that $\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)\geq 2\sqrt{2}$ and consequently for convergence with approximations, we obtain the following necessary condition on $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,$ from eq. 5,
$\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,<\frac{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}{2\sqrt{2}}.$ (6)
3.2 Error in subspace construction
In this section we consider the specific case of utilizing a lower precision to compute the matrix product $\textbf{H}\textbf{Y}^{(i)}_{k}$ in conjunction with approximating $\textbf{B}^{-1}$ with $\textbf{D}^{-1}$ in the Chebyshev recurrence relation defined by eq. 2. We first evaluate the upper bound on $\boldsymbol{\Delta}_{p}^{(i)}$ if one naively replaces the matrix product $\textbf{H}\textbf{Y}^{(i)}_{k}$ in eq. 2 with the approximate matrix product denoted as $\textbf{D}^{-1}\otimes\textbf{A}\otimes\textbf{Y}^{(i)}_{k}$}, where $\otimes$ represents the product evaluated with lower precision arithmetic, and argue that this method will fail to converge to the same residual tolerance that can be achieved with full precision matrix products and no approximation in computing the action of $\textbf{B}^{-1}$. We then propose a residual-based reformulation of the recurrence relation and argue that the proposed reformulation allows for convergence to similar residual tolerances that can be achieved with full precision matrix products and without approximations.
3.2.1 Traditional Chebyshev filtering method
Employing low-precision matrix-products and approximate inverse, the recurrence relation in eq. 2 for $\underline{\textbf{Y}}^{(i)}_{k}$, where the underline denotes that the matrix was constructed using approximations and $k=2,\dots,p$, can be written as
$\underline{\textbf{Y}}_{k+1}^{(i)}=a_{k}{\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}}+b_{k}\underline{\textbf{Y}}_{k}^{(i)}+c_{k}\underline{\textbf{Y}}_{k-1}^{(i)}$ (7)
where for convenience of notation we have defined
$a_{k}$ $=\frac{2\sigma_{k+1}}{e}$ $b_{k}$ $=-\frac{2\sigma_{k+1}c}{e}$ $c_{k}$ $=-\sigma_{k}\sigma_{k+1}$
and with the initial conditions $\underline{\textbf{Y}}_{0}^{(i)}=\textbf{X}^{(i)}$ and $\underline{\textbf{Y}}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}\textbf{A}-c\textbf{I})\textbf{X}^{(i)}$.
Theorem 3.2.
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using inexact matrix-products of the recurrence relation eq. 7 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ for $k=0,1,\dots,p$ where $\eta_{k}$ and $\tilde{\eta}_{k}$ are some finite constants that depend on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$
Proof.
This is proved in appendix B of the Appendix. ∎
We now discuss the implications of the theorem 3.2 within the context of eq. 5. Without loss of generality, we assume that our initial guess of trial subspace $\mathcal{S}^{(i)}$ at $i=0$ satisfies the inequality below as stated in eq. 5.
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\left(\sec{\angle(\mathcal{S}^{(i)},\mathcal{S})}+\csc{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (8)
While the above inequality is satisfied, $\angle(\mathcal{S}^{(i)},\mathcal{S})$ reduces as $i$ increases (from Theorem 3.1). However, we note that the RHS of this equation does not monotonically decrease with a decrease in $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and increases as $\angle(\mathcal{S}^{(i)},\mathcal{S})$ approaches 0. Thus, it stands to reason that beyond a certain value of $\angle(\mathcal{S}^{(i)},\mathcal{S})$, the inequality in eq. 8 no longer holds, and we cannot approach the wanted eigenspace beyond this point. Under the condition that $\gamma_{m}\eta_{p}+\zeta\tilde{\eta}_{p}<<\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$, we can estimate the closest angle that can be achieved and is given by the following expression
$\angle(\mathcal{S}^{(i)},\mathcal{S})\approx\frac{\gamma_{m}\eta_{p}+\zeta\tilde{\eta}_{p}}{\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,}$ (9)
We now propose a residual-based reformulation of ChFSI and argue that the proposed method does not suffer from this stagnation behavior even when employing lower precision arithmetic in matrix products.
3.2.2 Proposed residual-based Chebyshev filtering approach
To this end we define the weighted residual $\textbf{Z}^{(i)}_{k}$ in a given iteration $i$ for $k=0,\dots,p$ in the following way:
$\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)}))=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})\;\;\;\text{for}\;\;\;k=0,\dots,p$ (10)
where we have defined $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ and recall $p$ is the maximum Chebyshev polynomial degree used in the subspace construction step. The motivation behind this definition is the fact that both $C_{k}(\textbf{H})$ and H share the same eigenvectors with the eigenvalues being $C_{k}(\boldsymbol{\Lambda})$ and $\boldsymbol{\Lambda}$ respectively, this ensures that as $\textbf{X}^{(i)}$ and $\boldsymbol{\Lambda}^{(i)}$ approach the exact eigenvector and eigenvalues the weighted residual $\textbf{Z}_{k}^{(i)}$ approaches zero. Now, if we rewrite the recurrence relation defined in eq. 2 such that the matrix products are in terms of $\textbf{Z}_{k}^{(i)}$ then the relative error in the matrix products is then proportional to $\lVert\textbf{Z}_{k}^{(i)}\rVert\,$ thus ensuring a lower absolute error in the filtered subspace. We now propose the following recurrence relation that accomplishes our objective
Proposition 3.3.
The recurrence relation given by eq. 2 can be reformulated in terms of the weighted residuals defined by $\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}=\textbf{D}(\textbf{Y}_{k}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)})$ as
$\textbf{Z}_{k+1}^{(i)}$ $=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{Z}_{k}^{(i)}+b_{k}\textbf{Z}_{k}^{(i)}+c_{k}\textbf{Z}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$ (11)
$\boldsymbol{\Lambda}_{k+1}^{(i)}$ $=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$ (12)
where $\textbf{Z}_{0}^{(i)}=0$ and $\textbf{Z}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{D}\textbf{R}^{(i)}$ with $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$ and further we have $a_{k}=2\sigma_{k+1}/e$, $b_{k}=-2\sigma_{k+1}c/e$, and $c_{k}=-\sigma_{k}\sigma_{k+1}$.
Proof.
The recurrence relation for $\boldsymbol{\Lambda}_{k}^{(i)}=C_{k}(\boldsymbol{\Lambda}^{(i)})$ can be written as
$\boldsymbol{\Lambda}_{k+1}^{(i)}=a_{k}\boldsymbol{\Lambda}_{k}^{(i)}\boldsymbol{\Lambda}^{(i)}+b_{k}\boldsymbol{\Lambda}_{k}^{(i)}+c_{k}\boldsymbol{\Lambda}_{k-1}^{(i)}$
Multiplying with $\textbf{X}^{(i)}$, subtracting from eq. 2, and denoting $\textbf{Z}_{k}^{(i)}=\textbf{B}\textbf{R}_{k}^{(i)}$ we have
$\textbf{Z}_{k+1}^{(i)}$ $=a_{k}\textbf{D}\textbf{H}\textbf{D}^{-1}\textbf{Z}_{k}^{(i)}+b_{k}\textbf{Z}_{k}^{(i)}+c_{k}\textbf{Z}_{k-1}^{(i)}+a_{k}\textbf{D}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$
After computing $\textbf{Z}_{p}^{(i)}$ using this recurrence relation we can now evaluate $\textbf{Y}_{p}^{(i)}$ using the relation $\textbf{Y}_{p}^{(i)}=\textbf{D}^{-1}\textbf{Z}_{p}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{p}^{(i)}$. We note that using the approximation $\textbf{D}\textbf{B}^{-1}\approx\textbf{I}$ and employing lower precision matrix products in eq. 11, we can write the following recurrence relation
$\underline{\textbf{Z}}_{k+1}^{(i)}=a_{k}{\textbf{A}\otimes\textbf{D}^{-1}\otimes\underline{\textbf{Z}}_{k}^{(i)}}+b_{k}\underline{\textbf{Z}}_{k}^{(i)}+c_{k}\underline{\textbf{Z}}_{k-1}^{(i)}+a_{k}\textbf{B}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$ (13)
with $\textbf{Z}_{0}^{(i)}=0$ and $\textbf{Z}_{1}^{(i)}=\frac{\sigma_{1}}{e}\textbf{B}\textbf{R}^{(i)}$ and consequently, we have $\underline{\textbf{Y}}_{k}^{(i)}=\textbf{D}^{-1}\underline{\textbf{Z}}_{k}^{(i)}+\textbf{X}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}$. Note that this recurrence relation does not require the evaluation of $\textbf{B}^{-1}$.
Theorem 3.4.
The spectral norm of the error $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ in the subspace construction using recurrence relation given by eq. 13 satisfies $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq(\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k})\lVert\textbf{R}^{(i)}\rVert\,$ for $k=0,1,\dots,p$ where $\eta_{k}$ and $\tilde{\eta}_{k}$ are some finite constants that depends on $k$ and $\zeta=\lVert\textbf{D}^{-1}-\textbf{B}^{-1}\rVert\,$.
Proof.
This is proved in appendix C of the Appendix. ∎
Theorem 3.5.
The necessary condition for $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ for the case of residual-based Chebyshev filtering approach can be written as
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (14)
and if this inequality is satisfied for $i=i_{0}$ then it holds for all $i>i_{0}$
Proof.
From lemmas A.4, 3.4, and A.5 we have $\lVert\hat{\boldsymbol{\Delta}}_{p}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ and in order to have $\angle(\underline{\mathcal{S}}^{(i+1)},\mathcal{S})<\angle(\mathcal{S}^{(i)},\mathcal{S})$ from eq. 3 we require
$\lvert C_{p}(\lambda_{n})\rvert\,-\lvert C_{p}(\lambda_{n+1})\rvert\,$ $>2\lVert\textbf{H}\rVert\,\gamma_{m}\tilde{\eta}_{p}\left(1+\tan{\angle(\mathcal{S}^{(i)},\mathcal{S})}\right)$ (15)
We note that if this inequality is satisfied for some $i=i_{0}$, then it holds for all $i>i_{0}$ as the RHS decreases with decreasing $\angle(\mathcal{S}^{(i)},\mathcal{S})$ and hence we conclude that the R-ChFSI method can converge under approximations where the ChFSI method fails. ∎
Finally, the Chebyshev filtering step in R-ChFSI can be summarised in the following algorithm.
Algorithm 3 Residual based Chebyshev filtering procedure for generalized Hermitian eigenvalue problems
INPUTS: Chebyshev polynomial order $p$, estimates of the bounds of the eigenspectrum $\lambda_{max},\lambda_{min}$, estimate of the upper bound of the wanted spectrum $\lambda_{T}$ and the initial guess of eigenvectors $\textbf{X}^{(i)}$ and eigenvalues $\boldsymbol{\Lambda}^{(i)}$
OUTPUT: The filtered subspace $\textbf{Y}_{p}^{(i)}$
TEMPORARY VARIABLES:X, Y, $\textbf{R}_{\textbf{X}}$, $\textbf{R}_{\textbf{Y}}$$\boldsymbol{\Lambda}_{\textbf{X}}$ and $\boldsymbol{\Lambda}_{\textbf{Y}}$
$e\leftarrow\frac{\lambda_{max}-\lambda_{T}}{2}$; $c\leftarrow\frac{\lambda_{max}+\lambda_{T}}{2}$; $\sigma\leftarrow\frac{e}{\lambda_{min}-c}$; $\sigma_{1}\leftarrow\sigma$; $\gamma\leftarrow\frac{2}{\sigma_{1}}$
$\textbf{X}\leftarrow\textbf{X}^{(i)}$; $\textbf{Y}\leftarrow\textbf{A}\textbf{X}^{(i)}-\textbf{B}\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$
$\textbf{R}_{\textbf{X}}\leftarrow 0$; $\textbf{R}_{\textbf{Y}}\leftarrow\frac{\sigma_{1}}{e}\textbf{Y}$
$\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\textbf{I}$; $\boldsymbol{\Lambda}_{\textbf{Y}}\leftarrow\frac{\sigma_{1}}{e}\left(\boldsymbol{\Lambda}^{(i)}-c\textbf{I}\right)$
for $k\leftarrow 2$ to $p$ do
  $\sigma_{2}\leftarrow\frac{1}{\gamma-\sigma}$
  $\textbf{R}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}-\frac{2\sigma_{2}}{e}c\textbf{R}_{\textbf{Y}}-\sigma\sigma_{2}\textbf{R}_{\textbf{X}}+\frac{2\sigma_{2}}{e}\textbf{Y}\boldsymbol{\Lambda}_{\textbf{Y}}$
  $\boldsymbol{\Lambda}_{\textbf{X}}\leftarrow\frac{2\sigma_{2}}{e}\boldsymbol{\Lambda}_{\textbf{Y}}\boldsymbol{\Lambda}^{(i)}-\frac{2\sigma_{2}}{e}c\boldsymbol{\Lambda}_{\textbf{Y}}-\sigma\sigma_{2}\boldsymbol{\Lambda}_{\textbf{X}}$
  $\left(\textbf{R}_{\textbf{X}},\textbf{R}_{\textbf{Y}}\right); swap$$\left(\boldsymbol{\Lambda}_{\textbf{X}},\boldsymbol{\Lambda}_{\textbf{Y}}\right)$; $\sigma=\sigma_{2}$
end for
$\textbf{X}\leftarrow\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}+\textbf{X}\boldsymbol{\Lambda}_{\textbf{Y}}$
return X
We would also like to comment that while our above mathematical analysis is focused on generalized eigenvalue problems, it is trivial to propose a robust mixed precision eigensolver for standard eigenvalue problems by setting $\textbf{D}=\textbf{B}=\textbf{I}$. We note that the mixed precision R-ChFSI method for the standard eigenvalue problem is significantly more robust than the naive mixed-precision ChFSI algorithm, wherein we would perform the matrix products in eq. 2 in lower precision. The naive mixed precision ChFSI algorithm does not converge beyond a certain point, as evident from the discussion following eq. 8 and our numerical studies. In the interest of brevity, the subsequent section of this manuscript will focus solely on solutions for generalised eigenproblems using both low-precision and approximate inverses in computing matrix products for constructing the subspace rich in the desired subspace using Chebyshev filtering.
4 Results and Discussion
We present a detailed evaluation of the accuracy and efficiency achieved by the proposed residual-based Chebyshev filtered subspace iteration (R-ChFSI) method for solving sparse real symmetric and complex Hermitian generalized eigenvalue problems. We compare the performance of the R-ChFSI method (Algorithm 3) with that of traditional ChFSI approach (Algorithm 2), demonstrating its improved accuracy even when employing inexact matrix-vector products. Additionally, we investigate the computational performance of our proposed method on the Intel Data Centre GPU Max Series, utilising the Aurora supercomputing system, demonstrating its ability to maintain high accuracy while achieving greater computational efficiency.
For all our benchmark studies reported in this work, we consider the sparse matrix eigenvalue problems that arise from higher-order finite-element (FE) discretisation of Kohn-Sham density functional theory (DFT) [chenAdaptiveFiniteElement2014, schauerAllelectronKohnSham2013, tsuchidaAdaptiveFiniteelementMethod1996, tsuchidaElectronicstructureCalculationsBased1995, dasDFTFE10Massively2022, motamarriDFTFEMassivelyParallel2020, paskFiniteElementMethods2005, motamarriHigherorderAdaptiveFiniteelement2013], widely used in ab initio modelling of materials. The Kohn-Sham DFT equations, discretised using a non-orthogonal basis, are a canonical example of the large generalised Hermitian nonlinear eigenvalue problem [^3] The generalised eigenproblem arises in the finite-element discretisation of DFT equations because of the non-orthogonality of the FE basis functions. that motivates polynomial filtered-based methods. The nonlinear eigenvalue problem is solved as a sequence of linear eigenvalue problems using a self-consistent procedure (SCF). Each iteration of SCF requires a solution of the generalised Hermitian sparse eigenproblem for the lowest $n$ eigenpairs with $n<<m$, with $n$ proportional to the number of electrons in the material system and $m$ being the number of FE basis functions. Hence, we believe that Kohn-Sham DFT provides a demanding, realistic testbed for our algorithmic claims. For the sparse generalized eigenvalue problem of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$ considered here, we explore the use of diagonal approximation of B obtained by the mass lumping scheme, commonly employed in finite-element methods, to compute its inverse only during the subspace construction step within both the ChFSI and R-ChFSI methods, showing that the latter can achieve significantly lower residual tolerances.
We now evaluate the accuracy and performance of the proposed R-ChFSI method for solving the aforementioned Hermitian definite generalised eigenvalue problems arising in DFT by replacing algorithm 2 in algorithm 1 by algorithm 3, We implement the proposed R-ChFSI in the open-source code DFT-FE [^4] https://github.com/dftfeDevelopers/dftfe, a massively parallel finite-element code written in C++ for first principles based materials modelling using Kohn-Sham density functional theory (DFT). To this end, we consider a set of three benchmark material systems employing periodic boundary conditions, with the sparse matrix dimensions 1.728 million, 2.567 million, 85.766 million seeking 3000, 7000 and 13500 eigenpairs, respectively. These details are summarised in table 1. The three benchmark systems are chosen to be representative of a wide range of typical finite-element discretised DFT problems, considering the following factors: (a) differing spectral gap between wanted and unwanted eigenpsectrum, (b) differing number of target eigenpairs, and (c) both Gamma-point (real symmetric) and k-point (complex Hermitian) matrices. These distinctions matter for filtering-based eigensolvers. Systems with a small or zero spectral gap between wanted and unwanted spectra require higher amplification of the desired eigenspectrum and hence higher polynomial degree to separate near-degenerate states, while systems with a larger gap allow for lower polynomial degree and converge faster. Similarly, targeting a larger number of eigenvectors increases the cost of the Rayleigh-Ritz step (as the Rayleigh-Ritz step scales cubically, compared to the quadratic scaling of the filtering step) and alters the balance of where runtime is spent. Finally, k-point sampling of the Brillouin zone in DFT produces complex Hermitian matrices and different communication/compute patterns compared to Gamma-point runs, demonstrating that both cases ensure the conclusions apply to the typical varieties of DFT calculations encountered in practice.
Table 1: Dimensions of the benchmark problems considered. These correspond to 3 material systems comprising (1) $6\times 6\times 6$ supercell of Molybdenum, (2) $12\times 12\times 12$ supercell of Silicon and (3) $15\times 15\times 15$ supercell of Carbon, each with a single vacancy.
System # of DoFs (m) # of wanted eigenvectors (n) Subspace dimension
(1) 1728000 3000 3600
(2) 25672375 7000 8400
(3) 85766121 13500 16800
We will now demonstrate that the R-ChFSI method converges to a significantly lower residual tolerance than the ChFSI method when the diagonal approximation is employed for approximating the inverse overlap matrix during the subspace construction. To this end we will first consider the ChFSI and R-ChFSI algorithms for the benchmark systems summarized in table 1 using various Chebyshev polynomial filter degrees and compare the residual norm ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) that can be acheived for both the methods. We pick a target residual norm criteria as $\max_{j}r_{j}^{(i)}<10^{-8}$, a typical tolerance employed in DFT calculations. We subsequently proceed to analyse the behaviour of the R-ChFSI method when low-precision arithmetic is employed during the subspace construction. To this end, we benchmark the robustness and performance of the R-ChFSI algorithm when employing FP32 arithmetic. To this end, in algorithm 3 we store $\textbf{R}_{X}$ and $\textbf{R}_{\textbf{Y}}$ in FP32 and the computation of $\textbf{A}\textbf{D}^{-1}\textbf{R}_{\textbf{Y}}$ is done using FP32 arithmetic. Further, we also employ TF32 tensor cores (by using oneapi::mkl::blas::compute_mode::float_to_tf32), and we report the residual norms achieved by the FP32 and TF32 variants of R-ChFSI for our benchmark systems for various values of the Chebyshev polynomial degree. We further observe that using BF16 arithmetic does not provide any noticable performance during the filtering step and instead slows down the convergence, instead we chose to perform the computation using TF32 tensor cores and the nearest neighbour MPI communication required during FE discretized matrix multi-vector products in BF16 precision, we term this method TF32B for convenience of notation. This further improves the performance of the eigensolver, as will be demonstrated subsequently. For all benchmarking studies, we consider two cases: real symmetric and complex Hermitian eigenvalue problems.
4.1 Real Symmetric Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at the origin (Gamma-points) of the Brillouin zone for the systems described in table 1. This ensures that the resulting discretised equation is a real symmetric eigenvalue problem.
Figure 1: Plot of $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$ as the iterations progress for the ChFSI method and the R-ChFSI method (both in FP64 arithmetic) with various values of the Chebyshev polynomial filter degree ($p$ in Algorithm 2 and 3) to solve the symmetric generalized eigenvalue problem for the benchmark systems described in table 1.
Figure 2: Plot of $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$ as the iterations progress for the R-ChFSI method to solve the symmetric generalized eigenvalue problem with various precisions for the benchmark systems described in table 1. A slight offset has been added to the lower precision results for ease of visualization.
Figure 3: Speedups of lower precision R-ChFSI methods over the FP64 R-ChFSI method for subspace construction to solve the symmetric generalized eigenvalue problem for the benchmark systems described in table 1.
From fig. 1, we observe that for all the benchmark systems summarised in table 1 and various Chebyshev polynomial degrees, the residual tolerance that can be achieved by the R-ChFSI method is orders of magnitude lower than what can be achieved using the ChFSI method when employing FP64 arithmetic and the diagonal approximation for the inverse overlap matrix in the subspace filtering step. This observation is consistent with the analysis done in section 3. As previously discussed, the R-ChFSI algorithm enables us to construct the filtered subspace using lower precision arithmetic. We report the residual norms ($\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$) achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 3. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$ but not for $p=100,120$. We believe this is because the condition described in eq. 5 might not hold for higher values of $p$.
Figure 4: Speedups of lower precision R-ChFSI methods over the FP64 R-ChFSI method for the eigensolver to reach $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$ to solve the symmetric generalized eigenvalue problem for the benchmark systems described in table 1.
In fig. 3, we report the speedups achieved for the filtering step for Chebyshev polynomial degrees [^5] Polynomial degree 20 is not shown in fig. 3 as it did not converge to the desired tolerance within the time limit and degrees 100 and 120 are not shown as numerical instabilities start showing up at these polynomial degrees. of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. We achieve speedups of up to 1.8x for the TF32 variant and 2.6x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method. We further report in fig. 4 the speedups achieved for the total eigensolve that includes both filtering and Rayleigh-Ritz step (with the convergence criteria $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$) for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. For the full eigensolve, we achieve speedups of up to 1.6x for the TF32 and 2.0x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method.
4.2 Complex Hermitian Eigenvalue Problems
We now consider the finite-element discretization of the Kohn-Sham DFT equations sampled at a non-zero $k$-point  [dasDFTFE10Massively2022] in the Brillouin zone for the systems described in table 1. The resulting discretized equation is now a complex Hermitian eigenvalue problem.
Figure 5: Plot of $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$ as the iterations progress for the ChFSI method and the R-ChFSI method to solve the Hermitian generalized eigenvalue problem for the benchmark systems described in table 1.
Even for Hermitian eigenvalue problems, fig. 5 shows that across all benchmark systems summarized in table 1 and for the range of Chebyshev polynomial degrees, the R-ChFSI method attains residual tolerances that are orders of magnitude lower than those achievable with the standard ChFSI method when using FP64 arithmetic and the diagonal approximation for the inverse overlap matrix in the filtering step.
Figure 6: Plot of $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,$ as the iterations progress for the R-ChFSI method to solve the Hermitian generalized eigenvalue problem with various precisions for the benchmark systems described in table 1. A slight offset has been added to the lower precision results for ease of visualization.
Figure 7: Speedups of lower precision R-ChFSI methods over the FP64 R-ChFSI method for subspace construction to solve the Hermitian generalized eigenvalue problem for the benchmark systems described in table 1.
We also note that our proposed R-ChFSI algorithm allows us to construct the filtered subspace using lower precision for the complex Hermitian eigenproblem, and we report the residuals achieved by the FP32, TF32 and TF32B variants of R-ChFSI in fig. 6. We note that the residual norms obtained using the FP32, TF32 and TF32B variants are comparable to those obtained using the FP64 variant for Chebyshev polynomial degrees of $p=20,40,60,80$.
Figure 8: Speedups of lower precision R-ChFSI methods over the FP64 R-ChFSI method for the eigensolver to reach $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$ to solve the Hermitian generalized eigenvalue problem for the benchmark systems described in table 1.
We also report the speedups achieved for the filtering step for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 7 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. We achieve speedups of up to 2.3x for the TF32 variant and 2.7x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method. We further report the speedups achieved for the total eigensolve (with the convergence criteria $\max_{j}r_{j}^{(i)}=\max_{j}\lVert\textbf{A}\textbf{x}_{j}^{(i)}-\epsilon_{j}^{(i)}\textbf{B}\textbf{x}_{j}^{(i)}\rVert\,<10^{-8}$) for Chebyshev polynomial degrees of $p=40,60,80$ by the FP32, TF32 and TF32B variants of R-ChFSI over the FP64 variant of R-ChFSI in fig. 8 on Intel Data Center GPU Max Series accelerators deployed in the Aurora supercomputing system. For the full eigensolve, we achieve speedups of up to 1.9x for the TF32 and 2.1x for the TF32B variant of the R-ChFSI method over the FP64 variant of the R-ChFSI method.
5 Conclusion
In this work, we have proposed a residual-based Chebyshev filtered subspace iteration (R-ChFSI) method designed for large-scale sparse Hermitian eigenvalue problems, with a particular emphasis on robustness to inexact matrix–vector products. Through both mathematical analysis and extensive numerical experiments, we have demonstrated that the proposed method outperforms the standard Chebyshev filtered subspace iteration (ChFSI) approach significantly. Our theoretical justification demonstrates how reformulating the Chebyshev recurrence to operate on residuals, rather than directly on eigenvector estimates, effectively suppresses numerical errors introduced by inexact operator applications. This leads to a more reliable construction of the filtered subspace when compared with the traditional ChFSI recurrence relation.
An important contribution of this work lies in demonstrating the effectiveness of R-ChFSI for generalized Hermitian eigenvalue problems of the form $\textbf{A}\textbf{x}=\lambda\textbf{B}\textbf{x}$. In this setting, the proposed method is particularly impactful due to its ability to incorporate approximate inverses of B when constructing the subspace rich in the desired eigenspace using the Chebyshev polynomial recurrence relation. By replacing the exact computation of matrix inverses with computationally inexpensive approximate inverses (e.g., diagonal or block-diagonal approximations of B), the method not only reduces the overall computational burden but also maintains convergence robustness. This results in substantial performance gains, as the proposed approach achieves residual tolerances that are orders of magnitude lower than those obtainable with traditional ChFSI when solving the generalized problems. This highlights the method’s robustness and its suitability for large-scale discretised sparse generalized eigenproblems where exact factorisations of B are prohibitively expensive.
Another important aspect of the proposed approach is its natural compatibility with low-precision arithmetic. By design, R-ChFSI maintains stable convergence even when matrix–vector products are evaluated using lower-precision formats such as FP32 or TF32. Our results demonstrate that the method retains residual tolerances comparable to those obtained using full double precision, while benefiting from the reduced computational cost and increased throughput associated with low-precision operations. This makes R-ChFSI particularly attractive for emerging high-performance computing platforms, including GPU-accelerated systems and specialized tensor-processing hardware, which increasingly prioritize mixed-precision arithmetic to deliver high performance. The ability to jointly leverage approximate inverses and low-precision arithmetic allows R-ChFSI to achieve significant computational speedups while preserving numerical robustness.
The proposed R-ChFSI methodology has the potential to significantly impact a wide range of computational physics applications. These include ab initio modelling of materials, structural mechanics, fluid dynamics, and multiphysics problems, all of which rely on the efficient solution of large, sparse Hermitian eigenvalue problems arising from discretisations of partial differential equations using finite-element, finite-difference, wavelets or finite-volume formulations. By enabling the effective use of low-precision arithmetic without sacrificing accuracy, the method provides a compelling alternative for large-scale simulations that require substantial computational resources. Moreover, its robustness in solving large generalized nonlinear eigenvalue problems opens new opportunities for applying R-ChFSI to domains such as quantum modeling of materials, and other areas that demand scalable and reliable sparse eigensolvers.
6 Acknowledgements
The authors gratefully acknowledge the seed grant from the Indian Institute of Science and the SERB Startup Research Grant from the Department of Science and Technology, Government of India (Grant Number: SRG/2020/002194). This research used resources of the Argonne Leadership Computing Facility (ALCF), a US Department of Energy (DOE) Office of Science user facility at Argonne National Laboratory (ANL) and is based on research supported by the U.S DOE Office of Science-Advanced Scientific Computing Research Program, under Contract No. DE-AC02-06CH11357. PM acknowledges Vishwas Rao from ANL for useful discussions on the manuscript. The research also used the resources of PARAM Pravega at the Indian Institute of Science, supported by the National Supercomputing Mission (NSM), for testing a few of the benchmarks on CPU clusters. N.K. and K.R. would like to acknowledge the Prime Minister Research Fellowship (PMRF) from the Ministry of Education, India, for financial support. P.M. acknowledges the Google India Research Award 2023 for financial support during the course of this work.
7 Declaration of Generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used Paperpal in order to proofread. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Appendix
A Useful Lemmas
Lemma A.1.
Let the matrices $\textbf{E}\in\mathbb{C}^{m\times n}$ and $\textbf{E}_{\perp}\in\mathbb{C}^{m\times(m-n)}$ be given by $\textbf{E}=\Bigl[\begin{smallmatrix}\textbf{I}_{n}\\ 0\end{smallmatrix}\Bigr]$ and $\textbf{E}_{\perp}=\Bigl[\begin{smallmatrix}0\\ \textbf{I}_{m-n}\end{smallmatrix}\Bigl]$ where $\textbf{I}_{n}$ and $\textbf{I}_{m-n}$ are the $n\times n$ and $(m-n)\times(m-n)$ identity matrices respectively. Let $\mathcal{E}=\mathcal{R}(\textbf{E})$ and $\mathcal{E}_{\perp}=\mathcal{R}(\textbf{E}_{\perp})$ where $\mathcal{R}(.)$ denotes range of a matrix. Then for any matrix $\textbf{X}\in\mathbb{C}^{m\times n}$ with $\mathcal{X}=\mathcal{R}(\textbf{X})$ satisfying $\mathcal{E}_{\perp}\cap\mathcal{X}=\left\{0\right\}$ or equivalently $\text{rank}\left(\textbf{E}^{\dagger}\textbf{X}\right)=n$, we have
$\tan{\angle(\mathcal{X},\mathcal{E})}=\lVert\textbf{E}_{\perp}^{\dagger}\textbf{X}(\textbf{E}^{\dagger}\textbf{X})^{-1}\rVert\,$
Proof.
For proof of this statement, we refer the reader to [zhuAnglesSubspacesTheir2013, watkinsMatrixEigenvalueProblem2007]. ∎
Lemma A.2.
If $\textbf{H}\in\mathbb{C}^{m\times m}$ and $\textbf{X}\in\mathbb{C}^{m\times n}$, the spectral norm of the error in evaluating the matrix product HX due to floating point approximations satisfies $\lVert\textbf{H}\textbf{X}-{\textbf{H}\otimes\textbf{X}}\rVert\,\leq\gamma_{m}\lVert\textbf{H}\rVert\,\lVert\textbf{X}\rVert\,$
Proof.
This is a straightforward application of the results derived in [highamAccuracyStabilityNumerical2002, section 3.5]. We note that $\gamma_{m}$ is a constant that depends on $m$, and the machine precision of the floating point arithmetic (denoted as $\varepsilon_{M}$). For the case of dense real matrices, we have $\gamma_{m}=m^{2}\varepsilon_{M}/(1-m\varepsilon_{M})$ and for dense complex matrices, it needs to be appropriately modified as described in [highamAccuracyStabilityNumerical2002, section 3.6]. We also note that tigher bounds can be obtained from probabilistic error analysis [ipsenProbabilisticErrorAnalysis2020, highamSharperProbabilisticBackward2020]. Further, we also note that in the case where H is sparse, significantly tighter bounds independent of $m$ can be achieved depending on the specific implementations of the sparse matrix-dense matrix multiplication routines. ∎
Lemma A.3.
Consider a recurrence relation of the form
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$ (16)
with $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq h_{0}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+h_{1}^{(i)}\;\forall k=1,2,\dots,p$. We can then write
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+g_{k}^{1}h_{1}^{(i)}\right)$ (17)
Proof.
From eq. 16 we can write
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert b_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert c_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,+\lvert a_{k}\rvert\,\lVert\textbf{e}_{k}^{(i)}\rVert\,$
$\leq(\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+h_{0})\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lvert c_{k}\rvert\,\lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,+\lvert a_{k}\rvert\,h_{1}^{(i)}$
Defining the following
$\textbf{d}_{k}^{(i)}$ $=\begin{bmatrix}\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ \lVert\boldsymbol{\Delta}_{k-1}^{(i)}\rVert\,\end{bmatrix}$ $\textbf{F}_{k}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,\lVert\textbf{H}\rVert\,+\lvert b_{k}\rvert\,+\lvert a_{k}\rvert\,h_{0}&\lvert c_{k}\rvert\,\\ 1&0\end{bmatrix}$ $\textbf{E}_{k}^{(i)}$ $=\begin{bmatrix}\lvert a_{k}\rvert\,h_{1}^{(i)}\\ 0,\end{bmatrix}$
we can now write
$\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ $\leq\lVert\textbf{F}_{k}\textbf{d}_{k}^{(i)}+\textbf{E}_{k}^{(i)}\rVert\,\leq\lVert\textbf{F}_{k}\rVert\,\lVert\textbf{d}_{k}^{(i)}\rVert\,+\lvert a_{k}\rvert\,h_{1}^{(i)}$
Consequently, we now have
$\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ $\leq\left(\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,\right)\lVert\textbf{d}_{1}^{(i)}\rVert\,+\sum_{j=1}^{k}\left(\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\right)\lvert a_{j}\rvert\,h_{1}^{(i)}$
We note that $\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq\lVert\textbf{d}_{k+1}^{(i)}\rVert\,$ and $\lVert\textbf{d}_{1}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,$. This results in
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,\leq\left(\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,\right)\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+\sum_{j=1}^{k}\left(\prod_{r=j+1}^{k}\lVert\textbf{F}_{r}\rVert\,\right)\lvert a_{j}\rvert\,h_{1}^{(i)}$
Defining $g_{k}^{0}=\prod_{j=1}^{k}\lVert\textbf{F}_{j}\rVert\,$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\right)+\sum_{j=1}^{k}\frac{g_{k}^{0}}{g_{j}^{0}}\lvert a_{j}\rvert\,h_{1}^{(i)}$
$=g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+\sum_{j=1}^{k}\frac{\lvert a_{j}\rvert\,}{g_{j}^{0}}h_{1}^{(i)}\right)$
Defining $g_{k}^{1}=\sum_{j=1}^{k}\frac{\lvert a_{j}\rvert\,}{g_{j}^{0}}$ we have
$\lVert\boldsymbol{\Delta}_{k+1}^{(i)}\rVert\,$ $\leq g_{k}^{0}\left(\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,+\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,+g_{k}^{1}h_{1}^{(i)}\right)$
Lemma A.4.
The spectral norm of the residual, $\textbf{R}_{k}^{(i)}=C_{k}(\textbf{H})\textbf{X}^{(i)}-\textbf{X}^{(i)}C_{k}(\boldsymbol{\Lambda}^{(i)})$ for $k=0,\dots,p$, is bounded by $\lVert\textbf{R}_{k}^{(i)}\rVert\,\leq f_{k}\lVert\textbf{R}^{(i)}\rVert\,$ where $f_{k}$ is a finite constant and $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$
Proof.
Let $C_{k}(x)=\sum_{j=0}^{k}\alpha_{j}x^{j}$, we can now write
$\textbf{R}_{k}^{(i)}$ $=\sum_{j=1}^{k}\alpha_{j}\left(\textbf{H}^{j}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}^{j}\right)=\sum_{j=1}^{k}\alpha_{j}\sum_{r=0}^{j-1}\textbf{H}^{j-r-1}\left(\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\right){\boldsymbol{\Lambda}^{(i)}}^{r}$
Using submultiplicativity of the spectral norm we can now write
$\lVert\textbf{R}_{k}^{(i)}\rVert\,$ $\leq\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert\textbf{H}{\textbf{X}^{(i)}}-{\textbf{X}^{(i)}}{\boldsymbol{\Lambda}^{(i)}}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,=f_{k}\lVert\textbf{R}^{(i)}\rVert\,$
where we have defined $f_{k}=\sum_{j=0}^{k}\sum_{r=0}^{j-1}\lVert\alpha_{j}\textbf{H}^{j-r-1}\rVert\,\lVert{\boldsymbol{\Lambda}^{(i)}}^{r}\rVert\,$. ∎
Lemma A.5.
The norm of the residual $\textbf{R}^{(i)}$ satisfies $\lVert\textbf{R}^{(i)}\rVert\,\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$
Proof.
We have, $\textbf{R}^{(i)}=\textbf{H}\textbf{X}^{(i)}-\textbf{X}^{(i)}\boldsymbol{\Lambda}^{(i)}$ and the subspace diagonalization performed in the Rayleigh-Ritz projection step associated with $(i-1)^{th}$ iteration allows one to write $\boldsymbol{\Lambda}^{(i)}$ as $\boldsymbol{\Lambda}^{(i)}={\textbf{X}^{(i)}}^{\dagger}\textbf{H}\textbf{X}^{(i)}$. Subsequently we can write the residual $\textbf{R}^{(i)}=(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}$. We note that $\textbf{H}=\hat{\textbf{U}}_{1}\boldsymbol{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\boldsymbol{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger}$ and consequently we can write
$\lVert\textbf{R}^{(i)}\rVert\,=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\textbf{H}\textbf{X}^{(i)}\rVert\,=\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})(\hat{\textbf{U}}_{1}\boldsymbol{\Lambda}_{1}\hat{\textbf{U}}_{1}^{\dagger}+\hat{\textbf{U}}_{2}\boldsymbol{\Lambda}_{2}\hat{\textbf{U}}_{2}^{\dagger})\textbf{X}^{(i)}\rVert\,\\ \leq\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,\lVert\boldsymbol{\Lambda}_{1}\rVert\,\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,+\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{2}\rVert\,\lVert\boldsymbol{\Lambda}_{2}\rVert\,\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,$
Using the fact that $\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}={\lVert(\textbf{I}-\textbf{X}^{(i)}{\textbf{X}^{(i)}}^{\dagger})\hat{\textbf{U}}_{1}\rVert\,=\lVert\hat{\textbf{U}}_{2}^{\dagger}\textbf{X}^{(i)}\rVert\,}$ and the inequalities $\lVert\hat{\textbf{U}}_{1}^{\dagger}\textbf{X}^{(i)}\rVert\,\leq 1$, $\lVert(\textbf{I}-{\textbf{X}^{(i)}}^{\dagger}\textbf{X}^{(i)})\hat{\textbf{U}}_{2}\rVert\,\leq 1$ we can write
$\lVert\textbf{R}^{(i)}\rVert\,$ $\leq\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}(\lVert\boldsymbol{\Lambda}_{1}\rVert\,+\lVert\boldsymbol{\Lambda}_{2}\rVert\,)\leq 2\lVert\textbf{H}\rVert\,\sin{\angle(\mathcal{S}^{(i)},\mathcal{S})}$ (18)
B Proof of Theorem 3.2
Using the exact recurrence relation eq. 2 and the recurrence relation employing inexact matrix products eq. 7 we can now write a recurrence relation for $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}$ as
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}+a_{k}\textbf{e}_{k}^{(i)}$ (19)
with the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}$, obtained from the initial conditions of eq. 2 and eq. 7. While $\textbf{e}_{k}^{(i)}=\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}$. From lemma A.2 we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,=\lVert{\textbf{D}^{-1}\otimes\textbf{A}\otimes\underline{\textbf{Y}}_{k}^{(i)}}-\textbf{H}\underline{\textbf{Y}}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,$. We further note that
$\lVert\underline{\textbf{Y}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert\textbf{Y}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\textbf{X}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+\lVert C_{k}(\textbf{H})\rVert\,$
Thus we have $\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$. Using lemma A.3 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,=\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{X}^{(i)}\rVert\,$, $h_{0}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=(2\gamma_{m}\lVert\textbf{D}\rVert\,^{-1}+\zeta)\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k}$ with $\eta_{k}=2\lVert\textbf{D}^{(-1)}\rVert\,\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}$ and $\tilde{\eta}_{k}=\lVert\textbf{A}\rVert\,\lVert C_{k}(\textbf{H})\rVert\,g_{k-1}^{1}+g_{k-1}^{0}\lvert\frac{\sigma_{1}}{e}\rvert\,$.
C Proof of Theorem 3.4
Using the exact reformulated recurrence relation eq. 11 and the reformulated recurrence relation employing inexact matrix products eq. 13 we can now write a recurrence relation for the error as $\boldsymbol{\Delta}_{k}^{(i)}=\underline{\textbf{Y}}_{k}^{(i)}-\textbf{Y}_{k}^{(i)}=\textbf{D}^{-1}(\underline{\textbf{Z}}_{k}^{(i)}-\textbf{Z}_{k}^{(i)})$ as
$\boldsymbol{\Delta}_{k+1}^{(i)}=a_{k}\textbf{H}\boldsymbol{\Delta}_{k}^{(i)}+b_{k}\boldsymbol{\Delta}_{k}^{(i)}+c_{k}\boldsymbol{\Delta}_{k-1}^{(i)}\!+\!a_{k}\textbf{e}_{k}^{(i)}$
with $\textbf{e}_{k}^{(i)}=\left(\textbf{D}^{-1}-\textbf{B}^{-1}\right)\textbf{B}\textbf{R}^{(i)}\boldsymbol{\Lambda}_{k}^{(i)}+\textbf{D}^{-1}\textbf{A}\otimes\textbf{D}^{-1}\otimes\underline{\textbf{Z}}_{k}^{(i)}-\textbf{B}^{-1}\textbf{A}\textbf{D}^{-1}\underline{\textbf{Z}}_{k}^{(i)}$ and the initial conditions $\boldsymbol{\Delta}_{0}^{(i)}=0$ and $\boldsymbol{\Delta}_{1}^{(i)}=\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}$. We note that this recurrence relation is of the same form as that of eq. 16 in lemma A.3 with
$\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq\zeta(\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,+\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\underline{\textbf{Z}}_{k}^{(i)}\rVert\,)+2\gamma_{m}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,\lVert\textbf{D}^{-1}\underline{\textbf{Z}}_{k}^{(i)}\rVert\,$
We further note that using lemma A.4 we can write
$\lVert\textbf{D}^{-1}\underline{\textbf{Z}}_{k}^{(i)}\rVert\,=\lVert\boldsymbol{\Delta}_{k}^{(i)}+\textbf{D}^{-1}\textbf{Z}_{k}^{(i)}\rVert\,\leq\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,+f_{k}\lVert\textbf{D}^{-1}\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$
Thus we have
$\lVert\textbf{e}_{k}^{(i)}\rVert\,\leq(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\\ +((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)\lVert\textbf{R}^{(i)}\rVert\,$
Now, using lemma A.3 with $\lVert\boldsymbol{\Delta}_{0}^{(i)}\rVert\,=0$ and $\lVert\boldsymbol{\Delta}_{1}^{(i)}\rVert\,\!\!=\!\lVert\frac{\sigma_{1}}{e}(\textbf{D}^{-1}-\textbf{B}^{-1})\textbf{B}\textbf{R}^{(i)}\rVert\,\!\!\leq\lvert\frac{\sigma_{1}}{e}\rvert\,\zeta\lVert\textbf{B}\rVert\,\lVert\textbf{R}^{(i)}\rVert\,$, $h_{0}=(\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)\lVert\textbf{A}\rVert\,$ and $h_{1}^{(i)}=((\zeta+2\gamma_{m}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,)f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\zeta\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)\lVert\textbf{R}^{(i)}\rVert\,$ we obtain $\lVert\boldsymbol{\Delta}_{k}^{(i)}\rVert\,\leq(\gamma_{m}\eta_{k}+\zeta\tilde{\eta}_{k})\lVert\textbf{R}^{(i)}\rVert\,$ with $\eta_{k}=2f_{k}\lVert\textbf{D}^{-1}\rVert\,^{2}\lVert\textbf{D}\rVert\,\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,g_{k-1}^{1}g_{k-1}^{0}$ and $\tilde{\eta}_{k}=\lvert\frac{\sigma_{1}}{e}\rvert\,\lVert\textbf{B}\rVert\,g_{k-1}^{0}+(f_{k}\lVert\textbf{A}\rVert\,\lVert\textbf{D}^{-1}\textbf{B}\rVert\,+\lVert\textbf{B}\rVert\,\lVert\boldsymbol{\Lambda}_{k}^{(i)}\rVert\,)g_{k-1}^{1}g_{k-1}^{0}$.